OpenStudy (anonymous):

A circle with area A = pi r^2 is expanding with time. If dA/dt = 5cm^2/s, what is dr/dt when r=6?

OpenStudy (tkhunny):

Have you considered the 1st derivative in its differential form?

OpenStudy (anonymous):

well I am asking because I don't know what to do?

OpenStudy (tkhunny):

Non-responsive. You have \(A(r) = \pi r^{2}\). Can you find, from this definition, \(\dfrac{dA}{dr}\), the 1st derivative of A with respect to r?

OpenStudy (anonymous):

is the derivative 2r?

OpenStudy (tkhunny):

Excellent. Put the \(\pi\) in there and you'll have it.

OpenStudy (anonymous):

2pi r

OpenStudy (anonymous):

or is it 2r pi?

OpenStudy (tkhunny):

*** Side Bar *** \(A(r) = \pi r^{2}\) \(\dfrac{dA}{dr} = 2\pi r\) Did we just say the derivative of Area is the Circumference? I think we did! *** End of Side Bar *** Just for reference, it will help you if you try to use full and correct notation. You have the derivative: \(\dfrac{dA}{dt} = 2\pi r\) This is the 1st Derivative in its "Functional Notation". Now, we do a little magic trick. It's kind of like multiplying by "dr", but it really isn't. It's just a different way to write it. \(dA = 2\pi r\;dr\). This is still the first derivative. It's just the "Differential Form". In this form, we just have a little algebra problem. Can you identify the required pieces? \(dA = \) ?? \(dr = \) ?? \(r = \) ??

OpenStudy (tkhunny):

Multiplication is commutative. \(2\pi r\;and\;2r\pi\) are the same.

OpenStudy (anonymous):

the first thing I want to do is divide by dr

OpenStudy (tkhunny):

Why would you do that? The problem statement asks you to find dr. dA/dt = 5cm^2/s r=6 \(\dfrac{dA}{dt} = 2\pi r \dfrac{dr}{dt}\) Thus, \(5 = 2\pi(6) \dfrac{dr}{dt}\) -- Solve for dt (or dr/dt, however you write it.)